OpenStudy (anonymous):

2 over y+3 = 2 over 2y-2

OpenStudy (angela210793):

cross multiply 2(y+3)=2(2y-2) 2y+6=4y-4 -2y=-10 2y=10 y=10/2=5

OpenStudy (anonymous):

thank you (:

OpenStudy (angela210793):

:) uw